OpenStudy (anonymous):

solve. and get a medal nd fan ^2(√x-97)+1=(23)^2

OpenStudy (mathmale):

It might be worth your while to learn to use Equation Editor. I've had to guess what your equation represents. Is it\[\sqrt[2]{x-97)} + 1 = 23^2?\]